# Robi Butler: Remote Multimodal Interactions with Household Robot Assistant

### v1

@@ V-B1 Experimental Design
We recruited twelve volunteers (P1–P12; 7 males, 5 females from the university community. None of the participants had familiar experience with AR/MR equipment. We compared the performance of Robi Butler with two baseline systems by removing user interaction modalities, similar to an ablation study, resulting in three systems: Gesture-only, Voice-only, and Robi Butler (Gesture+Voice). In the Gesture-only system, buttons were added for participants to select the action to be executed. For the Voice-only system, we adapted the interactive visual grounding model from [48]. The use of the two baseline can be found in the websitefootnotemark: . Three representative tasks, T1 (object rearrangement), T2 (monitoring), and T3 (object rearrangement + monitoring), were selected from the previous experiment (Sec V-A1). As shown in Fig. 11, these tasks engaged the main areas of the home environment.
@@ V-B2 Analysis and Results
Fig. 10 shows the task performance of the three systems. A one-way repeated measures ANOVA was conducted to analyze the quantitative data after confirming normality assumptions. Both the Gesture-only and Gesture+Voice (i.e., Robi Butler) systems achieved a perfect task success rate of 100%, while the voice system had a slightly lower, though non-significant, success rate of 94.4%. This difference was attributed to errors in target referencing with voice commands only. For example, the voice recognition system misinterpreted the word ‘right’ as ‘red’, leading to the grounding error. Additionally, Robi Butler ($M=143.8$, $SD=14.8$) had a significantly lower task completion time than the Gesture-only system ($M=170.00$, $SD=21.4$) ($p<0.05$), but was not significantly lower than the Voice-only system ($M=157.1$, $SD=26.6$). The reduced task completion time for voice-supported systems was primarily due to the ability to use voice commands to express combined queries, whereas with the Gesture-only system, participants had to perform multiple manual button clicks, increasing task completion time.
Regarding the trust, the Robi Butler ($M=5.77$, $SD=0.97$) was perceived as significantly more trustworthy compared to both the Gesture-only system ($M=4.98$, $SD=0.85$, $p<0.05$) and the Voice-only system ($M=4.71$, $SD=1.03$, $p<0.05$). This suggests that combining gestures with voice commands fosters greater confidence in system reliability and consistency, outperforming systems relying solely on a single modality. P2 reasoned that “I trusted the gesture plus voice system the most because I found it easier to avoid making mistakes with it. For language only, sometimes it may misunderstand me. For gestures, I have to do the interaction multiple times.
For the SUS, participants gave the Gesture-only the lowest usability score ($M=47.29$, $SD=15.90$), which significantly lower than both Voice-only ($M=70.42$, $SD=11.62$, $p<0.01$) and Robi Butler ($M=75.83$, $SD=9.61$, $p<0.01$). This also indicates that Robi Butler achieved ‘Good’ usability (i.e., $SUS>75$  [52]) compared to the other systems.
Overall, the Robi Butler achieves the best performance, the highest usability, and the minimum perceived cognitive load among the baselines, answering RQ2. This was primarily due to the complementary nature of voice and gesture interactions, where voice enabled natural and combined queries. In contrast, gestures facilitated the disambiguation of voice commands related to locations and provided precise spatial annotations. Although multimodal interaction generally outperformed unimodal interaction, P10 expressed a negative sentiment, stating, “Using both voice and gesture is [sometimes] hard, as I need to switch between two modalities. I prefer voice-only as I don’t need to move my arm physically.” Future improvements, such as incorporating eye gaze tracking to minimize hand interactions, could potentially reduce such physical workload.
@@ VI Conclusion
This work introduces an interactive robotic assistant for household tasks using multimodal interactions with remote users. We outline three core components of the robot butler system and demonstrate its effectiveness in assistive question-answering and object rearrangement. Experiments show Robi Butler grounds multimodal instructions with a high task success rate, reasonable time, and minimal interactions. Follow-up tests confirm that combining voice and gestures enhances usability and trust, and reduces cognitive load compared to unimodal systems. In future work, we aim to make Robi Butler more adaptable, capable of autonomous skill learning, personalized interactions, and handling complex tasks that may require tactile feedback [53].
@@ Acknowledgment
We would like to express our gratitude to Yuhong Deng and Hanbo Zhang from the National University of Singapore for their invaluable support in model deployment. We also thank Tongmiao Xu from Xi’an Jiaotong University for her efforts in data collection. Furthermore, we sincerely appreciate the volunteers who contributed their time to participate in the human-robot interaction experiments.

### v2

@@ V-B1 Experimental Design
We recruited twelve volunteers (P1–P12; 7 males, 5 females) from the university community with IRB approval. None had prior experience with AR/MR smart glasses. We compared the performance of Robi Butler with two baseline systems by removing user interaction modalities, similar to an ablation study, resulting in three systems: Gesture-only, Voice-only, and Robi Butler (Gesture+Voice). In the Gesture-only system, buttons were added for participants to select the action to be executed. For the Voice-only system, we adapted the interactive visual grounding model from [52]. Videos of two baseline systems are available on the website. Three representative tasks, T1 (object rearrangement), T2 (monitoring), and T3 (object rearrangement + monitoring), were selected from the previous experiment (Sec V-A1). As shown in Fig. 11, these tasks engaged the main areas of the home. Participants were instructed to use their preferred verbal expressions.
@@ V-B2 Analysis and Results
Fig. 10 shows the task performance of the three systems. A one-way repeated measures ANOVA was conducted to analyze the quantitative data after confirming normality assumptions. Both the Gesture-only and Gesture+Voice (i.e., Robi Butler) systems achieved a perfect task success rate of 100%, while the voice system had a slightly lower, though non-significant, success rate of 94.4%. This difference was attributed to errors in target referencing with voice commands only. For example, the voice recognition system misinterpreted the word ‘right’ as ‘red’, leading to the grounding error. Additionally, Robi Butler ($M=143.8$, $SD=14.8$) had a significantly lower task completion time than the Gesture-only system ($M=170.00$, $SD=21.4$) ($p<0.05$), but was not significantly lower than the Voice-only system ($M=157.1$, $SD=26.6$). The reduced task completion time for voice-supported systems primarily resulted from the ability to use voice commands to express combined queries, whereas with the Gesture-only system, participants had to perform multiple manual clicks, increasing task completion time.
Regarding the trust, the Robi Butler ($M=5.77$, $SD=0.97$) was perceived as significantly more trustworthy compared to both the Gesture-only system ($M=4.98$, $SD=0.85$, $p<0.05$) and the Voice-only system ($M=4.71$, $SD=1.03$, $p<0.05$). This suggests that combining gestures with voice enhances the system’s reliability and consistency, outperforming single-modality systems. P2 reasoned that “I trusted the gesture plus voice system the most because I found it easier to avoid making mistakes with it. For language only, sometimes it may misunderstand me. For gestures, I have to do the interaction multiple times.” For the SUS, participants gave the Gesture-only the lowest usability score ($M=47.29$, $SD=15.90$), which significantly lower than both Voice-only ($M=70.42$, $SD=11.62$, $p<0.01$) and Robi Butler ($M=75.83$, $SD=9.61$, $p<0.01$). This also indicates that Robi Butler achieved ‘Good’ usability (i.e., $SUS>75$  [56]) compared to the other systems.
Overall, the Robi Butler achieves the best performance, the highest usability, and the minimum perceived cognitive load among the baselines, answering RQ2. This was primarily due to the complementary nature of voice and gestures—voice enabled natural queries, while gestures provided precise spatial annotations. Although multimodal interaction generally outperformed unimodal interaction, P10 expressed a negative sentiment, stating, “Using both voice and gesture is [sometimes] hard, as I need to switch between two modalities. I prefer voice-only as I don’t need to move my arm physically.” Incorporating eye gaze tracking could reduce physical workload by minimizing hand interactions.
@@ VI Conclusion
This work introduces an interactive robotic assistant for household tasks using multimodal interactions with remote users. We outline three core components of the robot butler system and demonstrate its effectiveness in assistive question-answering and object rearrangement tasks. Experiments show Robi Butler grounds remote multimodal instructions with a high task success rate, reasonable time, and minimal interactions. Follow-up tests confirm that combining voice and gestures enhances usability and trust, and reduces cognitive load compared to unimodal systems. In future work, we aim to make Robi Butler more adaptable, capable of lifelong learning, personalized interactions, and handling complex tasks that may require tactile feedback [57].
@@ Acknowledgment
We would like to thank Yuhong Deng, Hanbo Zhang, and Tongmiao Xu for their invaluable assistance in model development and data collection. We also sincerely appreciate all the volunteers who gave their time to participate in the human-robot interaction experiments. This research is supported in part by the National Research Foundation (NRF), Singapore and DSO National Laboratories under the AI Singapore Program (AISG Award No: AISG2-RP-2020-016) and by the Agency for Science, Technology & Research (A*STAR), Singapore under its National Robotics Program (No. M23NBK0091).